OpenStudy (anonymous):

Find the cross product a × b. a = j + 8k, b = 4i − j + 2k

OpenStudy (anonymous):

\[ {\rm if:}\quad \vec{a}=a_1\hat{i}+a_2\hat{j}+a_3\hat{k}\\ {\rm and}\quad \vec{b}=b_1\hat{i}+b_2\hat{j}+b_3\hat{k}\\ {\rm then}\\ \vec{a}\times\vec{b}=\left|\begin{matrix}\hat{i}&\hat{j}&\hat{k}\\ a_1&a_2&a_3\\ b_1&b_2&b_3 \end{matrix}\right| \] this is the determinant. plug in the numbers

OpenStudy (linyu):

\[\left[\begin{matrix}i & j & k \\ 0 & 1 & 8\\ 4 & -1 & 2 \end{matrix}\right]=i \left[\begin{matrix}1 & 8 \\ -1 & 2\end{matrix}\right]-j \left[\begin{matrix}0 & 8 \\ 4 & 2\end{matrix}\right]+k \left[\begin{matrix}0 & 1 \\ 4 & -1\end{matrix}\right]\] \[=i[(1)(2)-(-1)(8)]-j[(0)(2)-(4)(8)]+k[(0)(-1)-(4)(1)]\]\[=10i+32j-4k\]

OpenStudy (anonymous):

@Linyu nice work. but you have to follow the code of conduct. http://openstudy.com/code-of-conduct Cannot give away the answers just like that

OpenStudy (linyu):

@electrokid I understand and apologize